# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v11

@@ 5.1 Estimation Strategy
[…]
@@
[…]
@@ 5.2 Main Results
Table 2: Coal Mining and Regional Development: Local Population Growth in the Mining Area
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Panels A and B present the results for the 1930 and 1935 samples, respectively. Columns 1–3 show the estimates based on the IV estimator for the system of equations 2 and 3. Column 4 shows the estimates from the OLS estimator for equation 2. Panel A of Online Appendix Table C.4 shows the summary statistics for the population (before log-transformed) by treatment and control groups.
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.353535The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions; this shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when I use a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator. Then I include the city and town fixed effects (column 2) and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when I include the fixed effects. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
@@ 5.3 Fertility Responses
Mean of the DV $8.59$ $34.06$ $67.67$ $9.04$ $34.22$ $67.92$
Std. Dev. of the DV $2.28$ $5.52$ $11.22$ $2.41$ $5.47$ $11.07$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Columns 1–3 and 4–6 present the results for the 1930 and 1935 samples, respectively. Crude marriage, crude fertility, and marital fertility indicate the crude marriage rate, crude birth rate, and marital fertility rate, respectively Panel A of Online Appendix Table C.4 shows the summary statistics on these variables by treatment and control groups.
@@
Column 1 indicates that the estimate for the crude marriage rate is negative and statistically significant. This may reflect that the influx of single male miners could be associated with lower marriage rates among the entire local population. The estimate for the crude birth rate is also negative and statistically significant. This is consistent with the lower marriage rate in the mining area and the theoretical prediction that a relatively higher wage rate for female miners before the enforcement of regulation decreases the fertility rate. To disentangle both pathways, I use the marital fertility rate as the dependent variable in column 3. Although the denominator of this rate includes single females, it can rule out the attenuation from the greater proportion of single males in the mining area. The estimated coefficient is still statistically significantly negative. The estimate ($-10.72$) is roughly one standard deviation of the dependent variable, suggesting the economically meaningful effect.
The results for the 1935 sample listed in columns 4–6 show a clear change after the enforcement of the female labor regulation. While column (1) shows that the estimate for the marriage rate is still statistically significantly negative, its magnitude is roughly half of the estimate for the 1930 sample. Similarly, the estimate for the crude fertility rate is roughly a quarter (column 2). Importantly, the estimate for marital fertility shows the largest magnitude reduction, approximately one-seventh (column 3). It follows that the estimates for the fertility rates are no longer statistically significant.
@@ 5.4 Mortality Changes
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Columns 1–2 and 3–4 show the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for the male death rate, whereas columns 2 and 4 show the results for the female death rate. Panel A of Online Appendix Table C.4 summarizes the summary statistics of the mortality rates by treatment and control groups.
@@
First of all, I illustrate a theoretical framework for mortality selection in utero to better interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.414141 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
Table 5: Testing the Occupational Hazards and Pollutionusing the Early-life Mortality Data
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Panels A and B show the results for the 1933 and 1938 Vital Statistics samples, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. Panel B of Table C.4 summarizes the summary statistics of the early-life mortality variables by treatment and control groups.
@@ 5.5 Social Changes
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Columns 1–2 and 3–4 present the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for sex ratio (male/female), whereas columns 2 and 4 show the results for average household size. Panel A of Online Appendix Table C.4 shows the summary statistics of these variables by treatment and control groups.
@@ A.1 Female Miners in Coal Mine
Figure A.1: Male (saki-yama) and Female Miners (ato-yama)
@@
Notes: Figure A.1a shows the skilled male miner (saki-yama) extracting coal and the female miner (ato-yama) supporting her husband at the face. Figure A.1b illustrates an ato-yama bringing coal using a bamboo basket (sura) at the gangway through the steep pits. Both figures were taken in the early Taishō period. Source: Fumoto (1961, p. 2).
@@
A significant event that changed the labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after World War I, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night, but the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. Table 1 shows that the number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were enacted in September. Mining companies responded to this decline by introducing coal conveyors (Tanaka and Oghino 1977, p. 79). Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. [^50] Since female miners carried out the coal selection process, females were still in danger of being exposed to the toxicities in the wastewater, even outside the pits. More than 15,000 women still worked in the mines in the late 1930s, which represented approximately 10% of all the miners measured.
A.2 Influence of Great Depression in Japan
@@
The impact of the Great Depression was not substantial in Japan compared to that in other countries (Miyamoto 2008, pp.  56–57; Abe 2008, p. 106; Blumenthal 1981, pp. 46). Blumenthal (1981, p.43) shows that the index of industrial production in Japan did not decrease dramatically between 1929 (=100) and 1930 (=97). In contrast, the US and the UK suffered much more significant reductions (81 in the US and 92 in the UK) in the same period. Blumenthal (1981, p. 45) also reports that the unemployment rates between 1930 and 1935 in the US, the UK, and Japan were approximately 27%, 15%, and 6%, respectively. Consequently, Japan’s GNP growth rate (from 1929 to 1930) was positive at $1.1$%, whereas it was $-7.7$ and $-0.1$, respectively, in the US and UK. This evidence supports the validity of using the systematic data from the census and vital statistics of the 1930s.
A.3 Shares of Coal and GSC Mines
@@
(a) Coal and oil mines
(b) Metal mines
Notes: This figure shows the number of coal and oil (Figure A.2a) and metal mines (Figure A.2b) documented in the AHCs of 1931 and 1936. GSC indicates gold, silver, and copper mines. Source: Created by the author. Data on mines are from the Association for Harmonious Cooperation (1932, 1937).
Figure A.2: The Number of Coal, Oil, GSC, and the Other Metal Mines
in 1931 and 1936
@@
Appendix Appendix B Analytical Framework Appendix
@@
Males and females derive utilities from consumption ($c$) and having children ($b$). For a spouse $j\in\{m,f\}$, additively separable utilitu function is specified as follows:
@@
$\footnotesize{\begin{split}h_{m}=\left(\frac{\kappa_{m}}{w_{m}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b}\end{split}}$ (12)
$\footnotesize{\begin{split}h_{f}=\left(\frac{\kappa_{f}}{w_{f}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b}.\end{split}}$ (13)
@@
Notes: Columns (a) to (c) are from the survey report on $88$ children using the nursery at Yamauchi Coal Mine in Chikuhō coalfield in 1915. The figures in parentheses are the median values. Column (d) shows the average number of operating days of coal mines in 1924, surveyed in the mining census. The average number of operating days is calculated as 365 minus the average number of days closed ($45.5$). Column (e) shows the utilization rate, which is defined as: $100\left(\sum_{c=1}^{88}n_{c}\right)/\left(\sum_{c=1}^{88}p_{c}\times(365-45.5)\right)$. I excluded several individuals whose statistics were unavailable or unreliable. Three individuals whose length of service was longer than the number of years they had used the nursery are excluded. Six individuals had more days in care than days in operation. These individuals are also excluded because of suspected listing and transcription errors. Finally, we excluded three individuals for whom statistics on the number of days of childcare were not available. As a result, this table uses statistics on the 88 individuals among $100$ children originally listed. Sources: Noyori 2010, pp. 80–82. The average number of days closed is obtained from the Statistics Bureau of the Cabinet 1926, p. 18.

### v12

@@ 5.1 Estimation Strategy
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+\mathbf{x}_{i}^{\prime}\bm{\gamma}+e_{i},\end{split}}$ (2)
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. Because the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. For instance, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.323232 For example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\gamma\textit{City}+e$. When the municipality type (City) is unobserved, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\gamma$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. Thus, when $y$ is the population, conditional on mine deposits, it is reasonable to suppose that the town has a greater number of people than villages ($\gamma>0$). This result implies that $\beta^{*}=\beta+\Xi\gamma>\beta$. To deal with this systematic bias in the ordinary least squares, I included two indicator variables that equal one for cities and towns, respectively. As explained in Section 4.4, I further include the distances to the nearest neighboring railway station, seaport, and river and the elevation to explicitly control for transportation accessibility. [^33] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, I can partially assess the randomness of the primary exposure variable (MineDeposit). As I show later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications, including the control variables, thereby supporting the randomness of the exposure variable. To be conservative, I prefer the specification including the control variables.
@@
$\footnotesize{\begin{split}\textit{MineDeposit}_{i}=\kappa+\tau\textit{Stratum}_{i}+\mathbf{x}^{\prime}_{i}\bm{\mu}+\epsilon_{i},\end{split}}$ (3)
@@ 5.2 Main Results
Table 2: Coal Mining and Regional Development:
Local Population Growth in the Mining Area
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B present the results for the 1930 and 1935 samples, respectively. Columns 1–3 show the estimates based on the IV estimator for the system of equations 2 and 3. Column 4 shows the estimates from the OLS estimator for equation 2. Panel A of Online Appendix Table C.4 shows the summary statistics for the population (before log-transformed) by treatment and control groups.
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.363636The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions; this shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when I use a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator. Then I include the city and town fixed effects (column 2) and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when I include the fixed effects. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
@@ 5.3 Fertility Responses
Mean of the DV $8.59$ $34.06$ $174.14$ $9.04$ $34.22$ $177.50$
Std. Dev. of the DV $2.28$ $5.52$ $29.69$ $2.41$ $5.47$ $29.75$
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–3 and 4–6 present the results for the 1930 and 1935 samples, respectively. Crude marriage, crude fertility, and marital fertility indicate the crude marriage rate, crude birth rate, and marital fertility rate, respectively Panel A of Online Appendix Table C.4 shows the summary statistics on these variables by treatment and control groups.
@@
Column 1 indicates that the estimate for the crude marriage rate is negative and statistically significant. This may reflect that the influx of single male miners could be associated with lower marriage rates among the entire local population. The estimate for the crude birth rate is also negative and statistically significant. This is consistent with the lower marriage rate in the mining area and the theoretical prediction that a relatively higher wage rate for female miners before the enforcement of regulation decreases the fertility rate. To disentangle both pathways, I use the marital fertility rate as the dependent variable in column 3, which can rule out the attenuation from the greater proportion of single males in the mining area. The estimated coefficient is statistically significantly negative. The estimate ($-36.79$) is more than one standard deviation of the dependent variable, suggesting the economically meaningful effect.
The results for the 1935 sample listed in columns 4–6 show a clear change after the enforcement of the female labor regulation. While column (1) shows that the estimate for the marriage rate is still statistically significantly negative, its magnitude is roughly half of the estimate for the 1930 sample. Similarly, the estimate for the crude fertility rate is roughly a quarter (column 2). Importantly, the estimate for marital fertility shows the largest magnitude reduction from the estimate in column 3. It follows that the estimates for the fertility rates are no longer statistically significant.
@@ 5.4 Mortality Changes
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–2 and 3–4 show the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for the male death rate, whereas columns 2 and 4 show the results for the female death rate. Panel A of Online Appendix Table C.4 summarizes the summary statistics of the mortality rates by treatment and control groups.
@@
First of all, I illustrate a theoretical framework for mortality selection in utero to better interpret the quantitative evidence from the vital statistics data. The biological sorting mechanism in utero suggested by the Trivers-Willard hypothesis (TWH) (Trivers and Willard 1973) implies that there are two possible cases in which fetuses are exposed to health shocks: scarring and selection.424242 Valente (2015) tests the TWH using the case of the civil conflict in Nepal and shows that fetal exposure to the war decreased the sex ratio at births.
@@
Table 5: Testing the Occupational Hazards and Pollution
using the Early-life Mortality Data
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B show the results for the 1933 and 1938 Vital Statistics samples, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. Panel B of Table C.4 summarizes the summary statistics of the early-life mortality variables by treatment and control groups.
@@ 5.5 Social Changes
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Columns 1–2 and 3–4 present the results for the 1930 and 1935 samples, respectively. Columns 1 and 3 show the results for sex ratio (male/female), whereas columns 2 and 4 show the results for average household size. Panel A of Online Appendix Table C.4 shows the summary statistics of these variables by treatment and control groups.
@@ A.1 Female Miners in Coal Mine
Notes: Figure A.1a shows the skilled male miner (saki-yama) extracting coal and the female miner (ato-yama) supporting her husband at the face. Figure A.1b illustrates an ato-yama bringing coal using a bamboo basket (sura) at the gangway through the steep pits. Both figures were taken in the early Taishō period.
Source: Fumoto (1961, p. 2).
@@
A significant event that changed the labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after World War I, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night, but the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. Table 1 shows that the number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were enacted in September. Mining companies responded to this decline by introducing coal conveyors (Tanaka and Oghino 1977, p. 79). [^51] See also Tanaka (1984, chapter. 4) for a comprehensive summary of the relationships between mechanization process and changes in labor patterns in the pits. Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. [^52] Since female miners carried out the coal selection process, females were still in danger of being exposed to the toxicities in the wastewater, even outside the pits. More than 15,000 women still worked in the mines in the late 1930s, which represented approximately 10% of all the miners measured.
A.2 Mechanization and Female Miners
The history illustrated in Online Appendix A.1 suggests that the female miners were more likely to be replaced via mechanization. If this anecdote captures the true story, the mines adopting many machines were less likely to rely on the female miners. To assess this point, I assess the relationship between the degree of mechanization and the use of female miners in coal mines. The systematic data capturing the mechanization in coal mines is unavailable. Fortunately, however, the Fukuoka Prefectural Statistics document the horsepower and the number of miners in the coal mines in Fukuoka, including the Chikuhō coalfield. Figure A.2 illustrates the raw relationship between the share of female miners and log-transformed horsepower in the 82 coal mines in Fukuoka in 1930. Overall, this confirms the negative correlation between both variables. The estimate from Tobit model dealing with several censored observations in Figure A.2 is $-4.197$, which is statistically significant at the $1$% level. This suggests that a one standard deviation increase in the log-transformed horsepower ($=2.0$) decreases the use of female mines by approximately $8.4$%, which accounts for more than half of the standard deviation of the share of female miners ($=15.6$). Thus, this result supports the idea that mechanization via the installation of new technologies has worked to replace female miners in the coal mines.
Figure A.2: Relationship between share of female miners (%) and log-horsepower (hp) in coal mines in Fukuoka
Note: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
Source: Created by the author. Data are obtained from the official report published by Fukuoka Prefecture (1932).
A.3 Influence of Great Depression in Japan
@@
The impact of the Great Depression was not substantial in Japan compared to that in other countries (Miyamoto 2008, pp.  56–57; Abe 2008, p. 106; Blumenthal 1981, pp. 46). Blumenthal (1981, p.43) shows that the index of industrial production in Japan did not decrease dramatically between 1929 (=100) and 1930 (=97). In contrast, the US and the UK suffered much more significant reductions (81 in the US and 92 in the UK) in the same period. Blumenthal (1981, p. 45) also reports that the unemployment rates between 1930 and 1935 in the US, the UK, and Japan were approximately 27%, 15%, and 6%, respectively. Consequently, Japan’s GNP growth rate (from 1929 to 1930) was positive at $1.1$%, whereas it was $-7.7$ and $-0.1$, respectively, in the US and UK. This evidence supports the validity of using the systematic data from the census and vital statistics of the 1930s. In the mining sector, while several mines in Chikuhō coalfield were documented to have ceased operations during the depression, these were temporary (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973). Figure 1 confirms that the reduction in coal production was marginal in the case of Japan. Overall, the depression mainly hit prices rather than production. While the price declines were relatively large in textiles, metals, and construction materials, however, the shocks influenced food and fuel prices less (Nakamura and Odaka 1989, p. 53).
A.4 Shares of Coal and GSC Mines
(a) Coal and oil mines
(b) Metal mines
@@ A.4 Shares of Coal and GSC Mines
Notes: This figure shows the number of coal and oil (Figure A.3a) and metal mines (Figure A.3b) documented in the AHCs of 1931 and 1936. GSC indicates gold, silver, and copper mines.
Source: Created by the author. Data on mines are from the Association for Harmonious Cooperation (1932, 1937).
@@
Appendix Appendix B Conceptual Framework Appendix
@@
Males and females derive utilities from consumption ($c$) and having children ($b$). For a spouse $g\in\{m,f\}$, additively separable utilitu function is specified as follows:
@@
$\footnotesize{\begin{split}h_{m}=\left(\frac{\kappa_{m}}{w_{m}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b},\end{split}}$ (12)
$\footnotesize{\begin{split}h_{f}=\left(\frac{\kappa_{f}}{w_{f}}\right)^{\frac{1}{\rho}}\left(\frac{1-\gamma}{\gamma}\right)^{\gamma}\{\kappa_{m}^{\frac{1}{\rho}}w_{m}^{\frac{1}{\rho-1}}+\kappa_{f}^{\frac{1}{\rho}}w_{f}^{\frac{1}{\rho-1}}\}^{\frac{\gamma\rho-1}{1-\rho}}y_{b},\end{split}}$ (13)
@@
Notes: Columns (a) to (c) are from the survey report on $88$ children using the nursery at Yamauchi Coal Mine in Chikuhō coalfield in 1915. The figures in parentheses are the median values. Column (d) shows the average number of operating days of coal mines in 1924, surveyed in the mining census. The average number of operating days is calculated as 365 minus the average number of days closed ($45.5$). Column (e) shows the utilization rate, which is defined as: $100\left(\sum_{c=1}^{88}n_{c}\right)/\left(\sum_{c=1}^{88}p_{c}\times(365-45.5)\right)$.
I excluded several individuals whose statistics were unavailable or unreliable. Three individuals whose length of service was longer than the number of years they had used the nursery are excluded. Six individuals had more days in care than days in operation. These individuals are also excluded because of suspected listing and transcription errors. Finally, we excluded three individuals for whom statistics on the number of days of childcare were not available. As a result, this table uses statistics on the 88 individuals among $100$ children originally listed. Sources: Noyori 2010, pp. 80–82. The average number of days closed is obtained from the Statistics Bureau of the Cabinet 1926, p. 18.